# Neuromodulation and homeostasis: complementary mechanisms for robust neural function

### v1

@@ RESULTS
Figure 3: Why do controlled neuromodulation and homeostasis lead to robust neural function? A. Trajectories of the STG population from Fig. 2 in the modulated conductance space during homogeneous scaling in tonic spiking only. As expected from homeostasis, all neurons move in the same direction. White circles represent the initial conditions at the start of the step, and the green triangle marks the final conditions. B. Same as panel A, but during the brief period when neuromodulation is active. All trajectories are parallel to one another. C. Same as panels A and B, but during the homeostasis phase following neuromodulation. The outcome depends on the type of neuromodulation: sharp (left) or controlled (right). Sharp neuromodulation results in non-robust outcomes, whereas controlled neuromodulation preserves robust neural function.
@@
In the first section, when the homeostasis controller was combined with sharp neuromodulation, unreliable outcomes were observed. To explain this behavior, we use a schematic representation of the modulated conductance plane (Fig. 4A). In this plane, different values of $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ can lead to similar mean intracellular calcium concentration. These values can be highlighted using calcium isoclines (green dashed lines in Fig. 4A). Likewise, different values of $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ can lead to similar firing activity. These values can be highlighted using activity isoclines (blue dashed lines in Fig. 4A). Calcium isoclines are parallel one to each other, whereas activity isoclines pass through the origin and rotate around this point. This behavior is explained by homeostatic tuning rules and homogeneous scaling, as different conductance ratios lead to different firing patterns. Initially, the neuron operates in the target spiking mode at the target calcium level, corresponding to the intersection of the spiking and the target calcium isoclines. When sharply neuromodulating the neuron, the modulated conductances change, and the neuron shifts to another state, now reaching a new target strong bursting isocline, which intersects a higher calcium isocline. Homeostasis then kicks in to reduce the calcium level to its target value. However, in the homeostatic controller, the “native” tonic spiking direction is encrypted, leading to changes in conductances along the spiking isocline. As a result, the neuron moves away from the bursting isocline, potentially disrupting the new function. This explains why sharp neuromodulation is incompatible with the homeostasis controller.
To harmonize the actions of both controllers and achieve robust functional modulation, we leverage the significant difference in their timescales. Neuromodulation controls the modulated ionic conductances, while homeostatic tuning rules regulate all ionic conductances over a slower timescale to maintain an appropriate calcium level. In the modulated conductance plane (Fig. 4B), the initial and neuromodulated states are the same as in Fig. 4A. However, in the case of controlled neuromodulation, both homeostatic and neuromodulatory controllers are active. Specifically, when homeostasis reduces the calcium level and shifts the neuron away from the bursting isocline, the neuromodulation action counteracts this perturbation in function by bringing the neuron back to the bursting isocline, preserving its function. The neuromodulatory controller operates several orders of magnitude faster than the homeostatic controller. As a result, fast corrections from the former render perturbations from the latter infinitesimal, leading to changes in the conductance ratio along the target activity isocline (strong bursting in this example). Ultimately, the neuron stabilizes at the intersection of the target calcium and target activity isoclines, where both controllers meet their objective.
@@
Figure 4: Schematic explanation of why controlled neuromodulation integrates better with homeostasis compared to sharp neuromodulation. A. Schematic of the modulated conductance space for homeostasis combined with sharp neuromodulation. Initially (white circles), the neuron spikes at the target calcium level, located at the intersection of the firing pattern isocline (dashed blue lines) and the calcium level isocline (dashed green lines). Following sharp neuromodulation to a new firing pattern (red diamond), calcium levels sharply increase. Homeostasis reduces calcium (solid red arrow) by moving along the same direction as before neuromodulation (dashed red arrow), causing the neuron to deviate from the neuromodulated firing pattern isocline. B. Same as panel A, but with controlled neuromodulation. The initial steps are identical, but after neuromodulation, as homeostasis acts to reduce calcium, controlled neuromodulation simultaneously adjusts to keep the neuron on the neuromodulated firing pattern isocline (sawtooth red arrow). Because controlled neuromodulation operates on a much faster timescale than homeostasis, the neuron remains on the neuromodulated firing pattern isocline throughout.
@@
By employing this tandem of controllers, neuronal function becomes robust to channel blockade, provided that the channel deletion is compensable. While homeostasis compensation alone might lead to unreliable recovery from channel blockade, incorporating controlled neuromodulation somehow stabilizes the neuron response to such blockade, even in the case of highly degenerate neuronal populations. Furthermore, if the channel blockade is physiologically recoverable—meaning the function of the blocked channel can be restored using other channels—there exists an optimal combination of neuronal feedback gains to preserve that function. In essence, if one channel is blocked, a specific concentration of particular neuromodulators can preserve the function.
For example, in Fig.5, various channel blockades are applied to a degenerate population of STG conductance-based models equipped with the combined homeostasis/neuromodulation tandem of controllers. The target behavior of the population is regular bursting. In Fig.5A, H-type channels are blocked. Here, the function is preserved even immediately after the blockade, without requiring compensation. Subsequently, the combined regulation of homeostasis and controlled neuromodulation maintains this function by restoring the neuron to an appropriate calcium level on a population-wide scale. In Fig. 5B and C, T-type calcium and calcium-controlled potassium channels are blocked, respectively. In these cases, the blockade results mostly in the loss of bursting function. However, by allowing the homeostasis/neuromodulation pair of controllers to compensate, the function is largely recovered, albeit with irregular bursting patterns in some neurons. Additionally, intracellular calcium levels stabilize at their target values. This preservation of functions is feasible because the actions of H-type, T-type calcium, and calcium-controlled potassium channels are degenerate and can be compensated for using other channels present within the cell. The STG neuron model incorporates numerous channels operating on slow and ultraslow timescales, enabling the compensation for lost channels in these timescales. It is worth noting that only the loss of the calcium-controlled potassium channel has been compensated using different neuronal feedback gain values, i.e., different concentrations of various neuromodulators, as in control conditions.
However, the blockade of fast transient sodium channels is irrecoverable (Fig. 5D), as these channels are crucial for initiating spikes on a fast timescale. Once the blockade is applied, compensation cannot restore spikes because the action of sodium channels is non-degenerate and occurs on the fast timescale of the neuron. Similarly, achieving the appropriate intracellular calcium level becomes unattainable. A comparable outcome would result from the blockade of delayed rectified potassium channels, which are responsible for the downstroke of the spikes.
@@
Figure 5: Controlled neuromodulation and homeostasis ensure the preservation of function under physiologically recoverable disturbances. A. Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during homeostasis with controlled neuromodulation (controlled changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ starting at the dashed blue line) for a degenerate population of 200 models with H-type ion channel blockade (left). The corresponding distribution of behaviors before the blockade, immediately after the blockade, and following compensation (right) demonstrates the effect of the blockade. In this case, no neurons lose bursting due to the channel blockade or compensation. B. Same as panel A, but with a CaT channel blockade. The blockade causes half of the neurons to lose bursting; however, all neurons recover bursting with compensation. C. Same as panel A, but with a KCa channel blockade. The blockade causes more than half of the neurons to lose bursting, but nearly all recover bursting with compensation. D. Same as panel A, but with a Na channel blockade. The blockade causes all neurons to lose bursting, and compensation does not restore bursting due to the non-degenerate essential role of Na channels.
@@
The application of this tandem of controllers extends even to networks, particularly central pattern generators (CPGs). CPGs are self-organized neuronal circuits mainly composed of bursting neurons connected primarily through inhibitory connections, capable of generating rhythmic output without external input (Yuste et al.,, 2005; Guertin,, 2009). One extensively studied circuit is responsible for the pyloric and gastric mill rhythms in crustaceans, observed in the stomatogastric ganglion. This network exhibits two rhythms at different speeds: the fast pyloric rhythm and the slow gastric mill rhythm. These rhythmic outputs are vital for coordinating muscle contractions in the stomach for digestion, among other functions. A minimalist network model of such a bi-rhythmic circuit is investigated in Fig. 6 (Gutierrez et al.,, 2013; Drion et al.,, 2019). It comprises two half-center oscillators operating at different speeds, connected via a central neuron. Half-center oscillators consist of two bursting neurons connected in mutual inhibition, enabling them to fire anti-phase bursts. In the network scheme, the purple and blue neurons constitute the fast half-center oscillator (representing the pyloric rhythm), while the green and yellow neurons form the slow half-center oscillator (representing the gastric mill rhythm). The red neuron serves as the central one, also in bursting mode. It has been demonstrated that the gastric mill rhythm requires a neuromodulatory input to be activated (Selverston et al.,, 2009).
In this study, the gastric mill rhythm can be reliably induced by the pair of controllers, but not with homeostasis combined with sharp neuromodulation. Under control conditions (Fig. 6A), i.e., without neuromodulatory input to the gastric mill rhythm, only the pyloric rhythm is active and the central neuron consistently bursts. In particular, variability in conductances is observed in this network (Fig. 6A right). When neurons are equipped with homeostasis only, activation of sharp neuromodulation for the gastric mill rhythm leads to rapid loss of function, due to the unreliability of this combination (Fig. 6B middle). However, when neurons undergo controlled neuromodulation, the gastric mill rhythm is successfully induced and sustained (Fig. 6B right). Despite the degeneracy of neurons, characterized by significant variability in conductance ratios, the network achieves rhythmic output with appropriate intracellular calcium levels. This toy experiment demonstrates that such a tandem of controllers can be applied in computational models to modulate and maintain not only single cell activity but also network-level activity.
@@
Figure 6: Central pattern generator networks can be robustly modulated to orchestrate their rhythmic output. A. Control condition of the simplified pyloric/gastric mill rhythm network (left), inspired from Gutierrez et al., (2013). Without neuromodulation, the fast rhythm is active (top two traces, middle), while the slow rhythm is inactive (bottom two traces, middle). In this case, all neurons in the network exhibit degenerate conductances (right). B. Schematic representation of the activated slow rhythm network (left), modulated by either sharp neuromodulation (middle) or controlled neuromodulation (right). With sharp neuromodulation, the slow rhythm fails due to its unreliability, whereas controlled neuromodulation sustains the slow rhythm.
@@ DISCUSSION
The efficient and robust integration of homeostasis and neuromodulation in neurons remains an open question with several proposed hypotheses (Marder et al.,, 2014; Marder and Bucher,, 2005; Cunha,, 2001). The improper integration of these mechanisms—for example, combining sharp neuromodulation with a homeostatic controller—can lead to negative interactions. Such interactions may result in unreliable behaviors or numerical instabilities.
@@
In contrast, pairing a homeostatic controller with controlled neuromodulation ensures reliable neural function modulation while maintaining appropriate calcium levels. This approach draws inspiration from biology, where neuromodulation is known to be activity-dependent, as observed experimentally (Kramer and Levitan,, 1990; Raymond et al.,, 1992). Specifically, the neuromodulation controller models the role of G-protein-coupled receptors and second messengers, embedding feedback in the metabolic neuromodulation cascade (Marcus and Bruchas,, 2023; Scheler,, 2004). This reliable interaction can be extended to neural circuits, such as the gastric mill circuit in crabs, to produce robust network-level functionality (Marder et al.,, 2014).
@@
However, the success of this interaction critically depends on a point in the conductance space where both controllers converge. This point represents a neural activity pattern that supports neuromodulated firing while maintaining homeostatic calcium levels. As shown in Fig. 4, this intersection of the strong bursting isocline (the target of controlled neuromodulation) and the target calcium level (the goal of homeostatic control) represents the convergence point. When such a point is absent—e.g., during sodium channel blockade (Fig. 5)—neither neuromodulated activity nor calcium homeostasis is sustained. This may lead to pathological compensations, where one controller prioritizes calcium regulation at the cost of abandoning the desired firing pattern (Starmer et al.,, 1991; Grant et al.,, 1984; Anger et al.,, 2001). Furthermore, discontinuities in the neuromodulated firing pattern may lead to transient pathological behaviors.
@@
Maximizing convergence between these control mechanisms requires increasing the likelihood of both an intersection and a continuous path in conductance space. We hypothesize that this can be achieved by maximizing neuronal degeneracy (Whitacre,, 2010; Rathour and Narayanan,, 2019; Whitacre and Bender,, 2010; Prinz,, 2017; Goaillard and Marder,, 2021). Higher degeneracy increases the likelihood that the functional and dynamical role of one ion channel can be compensated by others, thereby preserving both the path and the intersection. This underscores the essential role of degeneracy in enabling robust neural functions that can be modulated.
@@
Finally, this study underscores the potential risks associated with pharmacological ion channel blockade (X Cubeddu,, 2016; DeWitt and Waksman,, 2004; Imbrici et al.,, 2018). If neuronal degeneracy is insufficient, such interventions may disrupt neuron ability to maintain robust neuromodulation. Instead, we propose targeting elements within the neuromodulation cascade—such as second messengers or other components of the control loop—as more reliable pharmacological strategies. Such interventions would preserve the dynamics of controlled neuromodulation, providing a more robust alternative to direct ion channel blockade.
@@
The Julia programming language was used in this work Bezanson et al., (2017). Numerical integration was realized using DifferentialEquations.jl.
Conductance-based models
For all experiments, single-compartment conductance-based models were employed. These models articulate an ordinary differential equation for the membrane voltage $V$, where $N$ ion channels are characterized as nonlinear dynamic conductances, and the phospholipid bilayer is represented as a passive resistor-capacitance circuit. Mathematically, the voltage-current relationship of any conductance-based neuron model is expressed as follows:
@@
Here, $C$ represents the membrane capacitance, $g_{\mathrm{ion}}$ denotes the considered ion channel conductance and is non-negative, gated between 0 (all channels closed) and $\bar{g}_{\mathrm{ion}}$ (all channels open), $E_{\mathrm{ion}}$ and $E_{\mathrm{leak}}$ are the channel reversal potentials, $\mathcal{I}$ is the index set of intrinsic ionic currents considered in the model, and $I_{\mathrm{ext}}$ is the current externally applied in vitro, or the combination of synaptic currents. Each ion channel conductance is nonlinear and dynamic, represented by $g_{\mathrm{ion}}(V,t)=\bar{g}_{\mathrm{ion}}m_{\mathrm{ion}}^{a}(V,t)h_{\mathrm{ion}}^{b}(V,t)$, where $m_{\mathrm{ion}}$ and $h_{\mathrm{ion}}$ are variables gated between 0 and 1, modeling the opening and closing gates of ion channels, respectively. Throughout this study, both the isolated crab STG neuron model Liu et al., (1998) and the adapted DA neuron model Qian et al., (2014) (where SK channels had been blocked to enable bursting) were employed.
@@
The DA model consists of six ion channels that operate on various time scales: fast sodium channels ($\bar{g}_{\mathrm{Na}}$); delayed-rectifier potassium channels ($\bar{g}_{\mathrm{Kd}}$); L-type calcium channels ($\bar{g}_{\mathrm{CaL}}$); N-type calcium channels ($\bar{g}_{\mathrm{CaN}}$); ERG channels ($\bar{g}_{\mathrm{ERG}}$); and NMDA channels ($\bar{g}_{\mathrm{NMDA}}$). Note that, owing to the multicellular nature of NMDA channels, they were excluded from this study but were still used for simulations with baseline values.
@@
The homeostatic controller adjusts all conductances using an integration rule described in O’Leary et al., (2014). Let $\bar{g}_{i}$ denote the conductance of ion channel $i\in[1,N]$, and $m_{i}$ the corresponding mRNA concentration. The dynamics of the homeostatic controller are governed by the following equations:
@@
where $[Ca^{+2}]$ and $[Ca^{+2}]_{\mathrm{target}}$ represent the intracellular calcium concentration and its target value, respectively. In essence, the calcium concentration error is integrated into the mRNA levels, which subsequently modulate the corresponding channel conductances. This regulation aims to bring the calcium concentration to its target value, as higher channel conductances (particularly calcium-conducting ones) facilitate greater calcium influx. Starting from any initial condition, this controller drives the cell along a line in conductance space known as the homogeneous scaling line. This line passes through both the initial condition and the origin of the conductance space axes. Steady-state analysis reveals that the conductance ratios remain invariant, such that $\frac{\bar{g}_{i}}{\bar{g}_{j}}=\frac{\tau_{j}}{\tau_{i}}$. This invariance ensures that the relative scaling of conductances is preserved during homeostatic compensation. Consequently, the cell’s conductance properties develop along the homogeneous scaling beam defined by these ratios.
@@
The neuromodulation controller adjusts a subset of $n$ conductances, denoted as $\bar{g}_{\mathrm{mod}}\in\mathbb{R}^{n}$ with $n<N$, using the algorithm introduced in Fyon et al., (2023). This controller employs the concept of Dynamic Input Conductances (DICs) to determine target values for $\bar{g}_{\mathrm{mod}}$ based on the input neuromodulator concentration, represented as $\bar{g}_{0}\left([\mathrm{nmod}]\right)\in\mathbb{R}^{n}$. The error between the target and the current values, $e_{\mathrm{mod}}$, drives a Proportional-Integral (PI) controller, which updates $\bar{g}_{\mathrm{mod}}$ to track the desired reference. The controller dynamics are described by the following equations:
@@
where $K_{p}$ and $K_{i}$ are the proportional and integral gains of the PI controller, respectively. In summary, the controller adjusts the modulated conductances $\bar{g}_{\mathrm{mod}}$ in response to the local concentration of neuromodulator, producing new reference values for these conductances to fine-tune the cell’s firing pattern—effectively neuromodulating it. This neuromodulator controller produces changes in conductance ratios that are normally preserved within the homeostatic controller. The new reference values depend on the neuromodulator concentration, the type of neuron, and the cell’s current state, characterized by the full set of conductances $\bar{g}_{\mathrm{ion}}\in\mathbb{R}^{N}$. The computation of these references is achieved through DICs, which encapsulate this complexity.

### v2

@@ Results
A. Trajectories of the STG population from Fig 1 in the modulated conductance space during homogeneous scaling in tonic spiking only. As expected from homeostasis, all neurons move in the same direction. White circles represent the initial conditions at the start of the step, and the green triangle marks the final conditions.
B. Same as panel A, but during the brief period when neuromodulation is active. All trajectories are parallel to one another.
C. Same as panels A and B, but during the homeostasis phase following neuromodulation. The outcome depends on the type of neuromodulation: sharp (left) or controlled (right). Sharp neuromodulation results in non-robust outcomes, whereas controlled neuromodulation preserves robust neural function.
Under sharp neuromodulation, homeostasis operates alone in the bursting phase, following the scaling direction characteristic of its reference, i.e. unmodulated, state set by the relative time constants of mRNA transcription in the homeostasis model (Fig 3C, left). This results in a decrease in calcium levels, but as shown in Fig 1, the response is unreliable and often undesirable. Specifically, for most neurons, the homeostatic controller reduces $\bar{g}_{\mathrm{A}}$ toward negative values (saturated at 0 to prevent model instability) to compensate for $\bar{g}_{\mathrm{CaS}}$, driving the neuron away from robust bursting and leading to heterogeneous activity changes across the population.
In contrast, controlled neuromodulation allows the homeostatic and neuromodulation controllers to act together, maintaining the new conductance ratios associated with bursting (Fig 3C, right). As a result, homeostatic regulation follows a new homogeneous scaling trajectory aligned with bursting conductance ratios. These trajectories converge toward the origin, as in tonic spiking, but with a distinct direction determined by neuromodulation. Due to initial conductance degeneracy, the exact scaling varies across the population. The shift between the two homogeneous scalings – tonic spiking and bursting – emerges from neuromodulation, ensuring that the desired neuromodulated function is maintained.
@@
The incompatibility of sharp neuromodulation with homeostasis can be further analyzed using a schematic representation of the modulated conductance plane (Fig 4A). In this plane, different combinations of $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ can yield similar calcium levels, forming calcium isoclines (green dashed lines). Similarly, different conductance values can result in similar firing activity, forming activity isoclines (blue dashed lines). While calcium isoclines are parallel, activity isoclines radiate from the origin and rotate around it. This organization reflects homeostatic tuning rules: different conductance ratios correspond to distinct firing modes [38]. Initially, the neuron operates at the intersection of a target spiking isocline and a target calcium isocline. When sharp neuromodulation is applied, conductances shift to a new state corresponding to strong bursting, moving the neuron to a higher calcium isocline. Homeostasis then acts to restore calcium levels, but because it retains the ”native” tonic spiking scaling direction, it alters conductances along the spiking isocline rather than preserving the newly induced bursting state. As a result, the neuron drifts away from the bursting isocline, potentially disrupting its function. This explains why sharp neuromodulation leads to unreliable outcomes when coupled with homeostasis.
Fig 4: Schematic explanation of why controlled neuromodulation integrates better with homeostasis compared to sharp neuromodulation.
@@
A. Schematic of the modulated conductance space for homeostasis combined with sharp neuromodulation. Initially (white circles), the neuron spikes at the target calcium level, located at the intersection of the firing pattern isocline (dashed blue lines) and the calcium level isocline (dashed green lines). Following sharp neuromodulation to a new firing pattern (red diamond), calcium levels sharply increase. Homeostasis reduces calcium (solid red arrow) by moving along the same direction as before neuromodulation (dashed red arrow), causing the neuron to deviate from the neuromodulated firing pattern isocline.
B. Same as panel A, but with controlled neuromodulation. The initial steps are identical, but after neuromodulation, as homeostasis acts to reduce calcium, controlled neuromodulation simultaneously adjusts to keep the neuron on the neuromodulated firing pattern isocline (sawtooth red arrow). Because controlled neuromodulation operates on a much faster timescale than homeostasis, the neuron remains on the neuromodulated firing pattern isocline throughout.
To reconcile homeostasis and neuromodulation and achieve robust functional modulation, we leverage their distinct timescales. Neuromodulation controls specific ionic conductances, while homeostasis adjusts all conductances over a slower timescale to regulate calcium levels. In the modulated conductance plane (Fig 4B), the neuron initially follows the same trajectory as in Fig 4A. However, under controlled neuromodulation, both controllers remain active. When homeostasis lowers calcium levels and shifts the neuron away from the bursting isocline, the neuromodulatory controller counteracts this deviation, bringing the neuron back toward the bursting isocline and preserving the desired function. Because the neuromodulatory controller operates orders of magnitude faster than the homeostatic controller, its rapid corrections render homeostatic perturbations negligible. This interaction constrains changes in conductance ratios along the target activity isocline (strong bursting in this case). Ultimately, the neuron stabilizes at the intersection of the target calcium and target activity isoclines, where both controllers achieve their objectives.
@@
By employing this tandem of controllers, neuronal function becomes robust to channel blockade, provided that the channel deletion is compensable. While homeostasis alone may lead to unreliable recovery, incorporating controlled neuromodulation stabilizes the neuronal response to blockade, even in highly degenerate neuronal populations. Furthermore, if the blocked channel function can be physiologically recovered – meaning other channels can substitute for its role – there exists an optimal combination of neuronal feedback gains that preserves the function. In essence, when a channel is blocked, a specific concentration of neuromodulators can maintain the function.
For example, in Fig 5, various channel blockades are applied to a degenerate population of STG conductance-based models equipped with the combined homeostasis/neuromodulation tandem. The target behavior of the population is regular bursting. In Fig 5A, H-type channels are blocked. Here, the function remains intact even immediately after the blockade, without requiring compensation. The combined action of homeostasis and controlled neuromodulation then ensures that intracellular calcium levels are restored on a population-wide scale, maintaining the function.
Fig 5: Controlled neuromodulation and homeostasis ensure the preservation of function under physiologically recoverable disturbances.
@@
A. Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during homeostasis with controlled neuromodulation (controlled changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ starting at the dashed blue line) for a degenerate population of 200 models with H-type ion channel blockade (left). The corresponding distribution of behaviors before the blockade, immediately after the blockade, and following compensation (right) demonstrates the effect of the blockade. In this case, no neurons lose bursting due to the channel blockade or compensation.
B. Same as panel A, but with a CaT channel blockade. The blockade causes half of the neurons to lose bursting; however, all neurons recover bursting with compensation.
C. Same as panel A, but with a KCa channel blockade. The blockade causes more than half of the neurons to lose bursting, but nearly all recover bursting with compensation.
D. Same as panel A, but with a Na channel blockade. The blockade causes all neurons to lose bursting, and compensation does not restore bursting due to the non-degenerate essential role of Na channels.
In Fig 5B and C, T-type calcium and calcium-activated potassium channels are blocked, respectively. In these cases, the blockade largely disrupts bursting. However, allowing the homeostasis/neuromodulation controllers to compensate restores function in most neurons, albeit with some displaying irregular bursting. Additionally, intracellular calcium levels stabilize at their target values. This recovery is possible because the actions of H-type, T-type calcium, and calcium-activated potassium channels are degenerate, allowing their function to be compensated by other channels in the cell. The STG neuron model includes numerous channels operating on slow and ultraslow timescales, enabling compensation over these timescales. Notably, only the loss of the calcium-activated potassium channel has been compensated by adjusting neuronal feedback gains – equivalent to modifying neuromodulator concentrations – as in control conditions.
In contrast, the blockade of fast transient sodium channels is irrecoverable (Fig 5D), as these channels are essential for spike initiation on a fast timescale. Once the blockade is applied, compensation cannot restore spiking because sodium channel function is non-degenerate and operates on a timescale that alternative channels cannot compensate for. Consequently, intracellular calcium levels also fail to reach their target values. A similar outcome would occur following the blockade of delayed rectifier potassium channels, which are responsible for the downstroke of spikes.
@@
The application of this tandem of controllers extends beyond single neurons to networks, particularly central pattern generators (CPGs). CPGs are self-organized neuronal circuits composed primarily of bursting neurons connected through inhibitory synapses, capable of generating rhythmic output without external input [55, 17]. A well-studied example is the circuit responsible for the pyloric and gastric mill rhythms in crustaceans, located in the stomatogastric ganglion. This network produces two distinct rhythms: the fast pyloric rhythm and the slow gastric mill rhythm. These rhythmic outputs are essential for coordinating muscle contractions in the stomach for digestion, among other functions. A minimal network model of such a bi-rhythmic circuit is investigated in Fig 6 [18, 11]. It consists of two half-center oscillators operating at different speeds, connected via a central neuron. Half-center oscillators are formed by two bursting neurons connected through mutual inhibition, allowing them to fire anti-phase bursts. In this model, the purple and blue neurons constitute the fast half-center oscillator (representing the pyloric rhythm), while the green and yellow neurons form the slow half-center oscillator (representing the gastric mill rhythm). The red neuron acts as the central element, also in bursting mode. It has been shown that the gastric mill rhythm requires neuromodulatory input to be activated [47].
Fig 6: Central pattern generator networks can be robustly modulated to orchestrate their rhythmic output.
@@
A. Control condition of the simplified pyloric/gastric mill rhythm network (left), inspired from [18]. Without neuromodulation, the fast rhythm is active (top two traces, middle), while the slow rhythm is inactive (bottom two traces, middle). In this case, all neurons in the network exhibit degenerate conductances (right).
B. Schematic representation of the activated slow rhythm network (left), modulated by either sharp neuromodulation (middle) or controlled neuromodulation (right). With sharp neuromodulation, the slow rhythm fails due to its unreliability, whereas controlled neuromodulation sustains the slow rhythm.
In this study, the tandem of controllers reliably induces the gastric mill rhythm, whereas homeostasis combined with sharp neuromodulation does not. Under control conditions (Fig 6A), i.e., in the absence of neuromodulatory input to the gastric mill rhythm, only the pyloric rhythm remains active, with the central neuron bursting continuously. Notably, conductance variability is observed within the network (Fig 6A right). When neurons are equipped with homeostasis alone, the activation of sharp neuromodulation for the gastric mill rhythm leads to a rapid loss of function due to the unreliability of this combination (Fig 6B middle). However, with controlled neuromodulation, the gastric mill rhythm is successfully induced and sustained (Fig 6B right). Despite the intrinsic degeneracy of the neurons, reflected in significant variability in conductance ratios, the network achieves stable rhythmic output while maintaining appropriate intracellular calcium levels.
This toy experiment demonstrates that the tandem of homeostasis and controlled neuromodulation can be applied in computational models to modulate and sustain not only single-neuron activity but also network-level dynamics.
@@ Discussion
The efficient and robust integration of homeostasis and neuromodulation in neurons remains an open question, with several hypotheses proposed to explain their interaction [30, 6]. Improper integration of these mechanisms – for instance, combining sharp neuromodulation with a homeostatic controller – can lead to negative interactions, resulting in unreliable behaviors or numerical instabilities.
@@
In contrast, pairing a homeostatic controller with controlled neuromodulation enables reliable modulation of neural function while maintaining appropriate calcium levels. This approach is biologically inspired, as neuromodulation is known to be activity-dependent, as observed experimentally [24, 45]. Specifically, the neuromodulation controller models the role of G-protein-coupled receptors and second messengers, embedding feedback within the metabolic neuromodulation cascade [26, 46]. This reliable interaction extends to neural circuits, such as the gastric mill circuit in crabs, where it contributes to robust network-level function [30].
@@
However, the success of this interaction critically depends on a point in the conductance space where both controllers converge. This point represents a neural activity pattern that supports neuromodulated firing while maintaining homeostatic calcium levels. As shown in Fig 4, it corresponds to the intersection of the strong bursting isocline (the target of controlled neuromodulation) and the homeostatic calcium setpoint. When no such intersection exists – e.g., under sodium channel blockade (Fig 5) – neither neuromodulated activity nor calcium homeostasis can be sustained. This may lead to pathological compensations, where one controller prioritizes calcium regulation at the expense of the target firing pattern [48, 16, 1]. Additionally, discontinuities in neuromodulated firing patterns may induce transient pathological behaviors.
@@
Maximizing convergence between these control mechanisms requires increasing the probability of both an intersection and a continuous path in conductance space. We hypothesize that this can be achieved by maximizing neuronal degeneracy [54, 44, 53, 42, 15]. Higher degeneracy increases the likelihood that the functional and dynamical role of one ion channel can be compensated by others, thereby preserving both the path and the intersection. This highlights the crucial role of degeneracy in enabling robust, neuromodulated neural function.
@@
Finally, these findings underscore the potential risks associated with pharmacological ion channel blockade [5, 9, 21]. If neuronal degeneracy is insufficient, such interventions may compromise the ability to sustain robust neuromodulation. As an alternative, we propose targeting elements within the neuromodulation cascade – such as second messengers or other components of the control loop – as more reliable pharmacological strategies. By preserving the dynamics of controlled neuromodulation, these interventions could offer a more robust alternative to direct ion channel blockade.
@@
The Julia programming language was used in this work [3]. Numerical integration was realized using DifferentialEquations.jl.
Conductance-based model
For all experiments, single-compartment conductance-based models were employed. These models articulate an ordinary differential equation for the membrane voltage $V$, where $N$ ion channels are characterized as nonlinear dynamic conductances, and the phospholipid bilayer is represented as a passive resistor-capacitor circuit. Mathematically, the voltage-current relationship of any conductance-based neuron model is expressed as follows:
@@
Here, $C$ represents the membrane capacitance, $g_{\mathrm{ion}}$ denotes the considered ion channel conductance and is non-negative, gated between 0 (all channels closed) and $\bar{g}_{\mathrm{ion}}$ (all channels open), $E_{\mathrm{ion}}$ and $E_{\mathrm{leak}}$ are the channel reversal potentials, $\mathcal{I}$ is the index set of intrinsic ionic currents considered in the model, and $I_{\mathrm{ext}}$ is the current externally applied in vitro, or the combination of synaptic currents. Each ion channel conductance is nonlinear and dynamic, represented by $g_{\mathrm{ion}}(V,t)=\bar{g}_{\mathrm{ion}}m_{\mathrm{ion}}^{a}(V,t)h_{\mathrm{ion}}^{b}(V,t)$, where $m_{\mathrm{ion}}$ and $h_{\mathrm{ion}}$ are variables gated between 0 and 1, modeling the opening and closing gates of ion channels, respectively. Throughout this study, the isolated crab STG neuron model of [25] was employed.
@@
The homeostatic controller adjusts all conductances using an integration rule described in [38]. Let $\bar{g}_{i}$ denote the conductance of ion channel $i\in[1,N]$, and $m_{i}$ the corresponding mRNA concentration. The dynamics of the homeostatic controller are governed by the following equations:
@@
where $Ca^{+2}]$ and $Ca^{+2}]_{\mathrm{target}}$ represent the calcium level and its target value, respectively. In essence, the calcium level error is integrated into the mRNA levels, which subsequently modulate the corresponding channel conductances. This regulation aims to bring the calcium level to its target value, as higher channel conductances (particularly calcium-conducting ones) facilitate greater calcium influx. Starting from any initial condition, this controller drives the cell along a line in conductance space known as the homogeneous scaling line. This line passes through both the initial condition and the origin of the conductance space axes. Steady-state analysis reveals that the conductance ratios remain invariant, such that $\frac{\bar{g}_{i}}{\bar{g}_{j}}=\frac{\tau_{j}}{\tau_{i}}$. This invariance ensures that the relative scaling of conductances is preserved during homeostatic compensation. Consequently, the cell conductance properties develop along the homogeneous scaling beam defined by these ratios.
@@
The neuromodulation controller adjusts a subset of $n$ conductances, denoted as $\bar{g}_{\mathrm{mod}}\in\mathbb{R}^{n}$ with $n<N$, using the algorithm introduced in [14]. This controller employs the concept of Dynamic Input Conductances (DICs) to determine target values for $\bar{g}_{\mathrm{mod}}$ based on the input neuromodulator concentration, represented as $\bar{g}_{0}\left([\mathrm{nmod}]\right)\in\mathbb{R}^{n}$. The error between the target and the current values, $e_{\mathrm{mod}}$, drives a Proportional-Integral (PI) controller, which updates $\bar{g}_{\mathrm{mod}}$ to track the desired reference. The controller dynamics are described by the following equations:
@@
where $K_{p}$ and $K_{i}$ are the proportional and integral gains of the PI controller, respectively. In summary, the controller adjusts the modulated conductances $\bar{g}_{\mathrm{mod}}$ in response to the local concentration of neuromodulator, producing new reference values for these conductances to fine-tune the cell firing pattern – effectively neuromodulating it. This neuromodulator controller produces changes in conductance ratios that are normally preserved within the homeostatic controller. The new reference values depend on the neuromodulator concentration, the type of neuron, and the cell’s current state, characterized by the full set of conductances $\bar{g}_{\mathrm{ion}}\in\mathbb{R}^{N}$. The computation of these references is achieved through DICs, which encapsulate this complexity.